# Mpemba meets Newton: Exploring the Mpemba and Kovacs effects in the time-delayed cooling law

### v1

@@ Abstract
Despite extensive research, the fundamental physical mechanisms underlying the Mpemba effect, a phenomenon where a substance cools faster after initially being heated, remain elusive. Although historically linked with water, the Mpemba effect manifests across diverse systems, sparking heightened interest in Mpemba-like phenomena. Concurrently, the Kovacs effect, a memory phenomenon observed in materials like polymers, involves rapid quenching and subsequent temperature changes, resulting in nonmonotonic relaxation behavior. This paper probes the intricacies of the Mpemba and Kovacs effects within the framework of the time-delayed Newton’s law of cooling, recognized as a simplistic yet effective phenomenological model accommodating memory phenomena. This law allows for a nuanced comprehension of temperature variations, introducing a time delay ($\tau$) and incorporating specific protocols for the thermal bath temperature, contingent on a defined waiting time ($t_{\text{w}}$). Remarkably, the relevant parameter space is two-dimensional ($\tau$ and $t_{\text{w}}$), with bath temperatures exerting no influence on the presence or absence of the Mpemba effect or the relative strength of the Kovacs effect. The findings enhance our understanding of these memory phenomena, providing valuable insights applicable to researchers across diverse fields, ranging from physics to materials science.
@@ I Introduction
As a teenager, Erasto Bartholomeo Mpemba (1950–2023) accidentally discovered the paradoxical effect that now bears his name. In the first part of his renowned paper [1] with Denis Gordon Osborne (1932–2014), the 19-year-old Mpemba candidly recounted the story. In the second part of the paper, Osborne wrote “The headmaster of Mkwawa High School invited me to speak to the students on ‘Physics and national development’. […] One student raised a laugh from his colleagues with a question I remember as ‘If you take two beakers with equal volumes of water, one at $35^{\circ}$ and the other at $100^{\circ}$, and put them into a refrigerator, the one that started at $100^{\circ}$ freezes first. Why?’. It seemed an unlikely happening, but the student insisted that he was sure of the facts. I confess that I thought he was mistaken but fortunately remembered the need to encourage students to develop questioning and critical attitudes. No question should be ridiculed. In this case there was an added reason for caution, for everyday events are seldom as simple as they seem and it is dangerous to pass a superficial judgment on what can and cannot be. I said that the facts as they were given surprised me because they appeared to contradict the physics I knew. But I added that it was possible that the rate of cooling might be affected by some factor I had not considered. I promised I would put the claim to the test of experiment and encouraged my questioner to repeat the experiment himself.” In that second part of Ref. [1], Osborne reported some experimental results confirming the effect, although he conceded that “The experiments attempted were relatively crude and several factors could influence cooling rates. More sophisticated experiments are needed to provide a more certain answer to the question.”
@@
After 1969, attention to the Mpemba effect was mainly confined to popular science and education journals [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. It was also revealed that the phenomenon had already been noted by classical philosophers and scholars, including Aristotle[24], Roger Bacon [25], Francis Bacon [26], and René Descartes [27]. Interestingly, in the same year that Mpemba and Osborne published their paper, Kell independently published a brief note that began with the statement: “It is widely believed, at least in Canada, that hot water will freeze more quickly than cold water.”
@@
Figure 1: Number of publications and citations in the period 2006–2023 obtained from the Web of Science with the search query “topic=(Mpemba effect).” Note that the data for the year 2023 are not final.
In addition to its historical association with water, the Mpemba effect exhibits similar phenomena reported across a diverse range of systems, including carbon nanotube resonators [41], clathrate hydrates [42], Markovian models [43, 44, 45, 46, 47, 48], granular gases [49, 50, 51, 52, 53, 54, 55, 56, 57, 58], spin glasses [59], molecular gases under drag [60, 61, 62], non-Markovian mean-field systems [63, 64], inertial suspensions [65, 66], colloidal systems [67, 68, 45, 69, 70] quantum systems [71, 72, 73], Ising-like models [74, 75, 76, 48, 77], ideal gases [78], phase transitions [79], active matter [80], the autonomous information engine [81], ionic liquids [82], polymers [83, 84], or Langevin systems [85] The escalating interest in Mpemba-like effects is depicted in Fig. 1.
Figure 2: Sketch of supposed cooling curves for two similar systems in same environment. Adapted from Ref. [8].
@@
where $T_{\text{b}}$ is the temperature of the thermal bath (or environment) and $\lambda$ is the coefficient of heat transfer, here assumed to be a (positive) constant. The general solution is simply
@@
with $T_{0}=T(0)$ being the initial temperature. On the other hand, if the Mpemba effect exists in a certain material, memory dynamics must be taken into account, so that “the overtaking effect should enable us to describe the difference between the two systems as represented by the points B and D [see Fig. 2] in such a way that we would expect the system starting hotter to cool faster even below this temperature.” The simplest way of incorporating those memory phenomena into Newton’s law involves postulating that the temperature’s rate of change at time $t$ depends on the temperature at a preceding time $t-\tau$ [89], i.e.,
@@
where $\tau>0$ is the time delay and we have considered the possibility that the bath temperature, $T_{\text{b}}(t)$, changes with time. Equation (3) can be considered as the simplest phenomenological equation incorporating memory effects.
@@
Additionally to the Mpemba effect, another fascinating memory phenomenon is the Kovacs effect, originally reported in polymer materials [90, 91] and also observed in other complex systems [92, 93, 94, 95, 96, 97, 98, 99, 57]. In the Kovacs effect, a sample, initially in equilibrium at a high temperature $T_{\text{b}}^{-}$, undergoes rapid quenching to a lower temperature $T_{\text{b}}^{+}$. The sample evolves for a specified waiting time $t_{\text{w}}$, but subsequently the bath temperature is abruptly raised to $T(t_{\text{w}})$. Beyond $t=t_{\text{w}}$, the sample temperature exhibits a dynamic behavior: it initially increases, reaches a maximum, and then returns to equilibrium for a longer period. The effect thus highlights the nontrivial impact of the material’s thermal past on its present and future behavior.
@@
The aim of this paper is to analyze the Mpemba and Kovacs effects as described by the time-delayed cooling law, Eq. (3). Its general solution is studied in Sec. II, with special attention to single and double quenches. It is seen that, in the former case, the solution has a form similar to Eq. (2), except that the role of the exponential $e^{-\lambda t}$ is played by a function $\mathcal{E}(t)$, here called the $\tau$-exp function. This function is positive definite only if the time delay is smaller than a threshold value $\tau_{\max}=\lambda^{-1}e^{-1}$. Next, in Sec. III, the solution is applied to the study of the Mpemba effect under simple protocols involving a hot bath at temperature $T_{\text{b}}^{\text{h}}$ and a cold bath at temperature $T_{\text{b}}^{\text{c}}$. Sample A is thermalized at temperature $T_{\text{b}}^{\text{h}}$ and then quenched to temperature $T_{\text{b}}^{\text{c}}$. Conversely, sample B is thermalized at $T_{\text{b}}^{\text{c}}$, quenched to $T_{\text{b}}^{\text{h}}$, and, after a waiting time $t_{\text{w}}$, quenched again to the cold temperature $T_{\text{b}}^{\text{c}}$. For simplicity, the quench of sample A and the first quench of sample B occur at the same time. After this preparation protocol, both samples are in the same environment (bath $T_{\text{b}}^{\text{c}}$) but start with different temperatures. It is proved in Sec. III that (i) the existence or absence of the Mpemba effect is independent of the values of $T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$, (ii) the Mpemba effect exists if and only if the two control parameters ($\tau$ and $t_{\text{w}}$) lie inside a certain narrow region, and (iii) the direct (cooling process) and the inverse (heating process) effects are fully equivalent. Section IV is devoted to the Kovacs effect and the function characterizing the strength of the Kovacs hump is identified. Interestingly, this hump points downward in the cooling process, thus signaling the presence of an anomalous Kovacs effect [94, 99], which becomes relatively stronger as both the time delay $\tau$ and the waiting time $t_{\text{w}}$ increase. Finally, the paper is closed in Sec. V with a summary and conclusion.
@@ II Time-delayed Newton’s cooling law
II.1 General solution in Laplace space
The general solution of Eq. (3) depends not only on the initial temperature $T_{0}$ but it is actually a functional of the previous history $T(t)$ for $-\tau<t<0$. Given the linear character of Eq. (3), it is convenient to define the Laplace transforms
@@ II.1 General solution in Laplace space
and, for simplicity, we have taken $\lambda^{-1}=1$ as unit of time. Thus, the general solution in Laplace space is
@@
II.2 Basic solution. Single quench
@@
The system is assumed to have equilibrated at $T_{\text{b}}^{-}$ for $t<0$, that is, $T(t<0)=T_{\text{b}}^{-}$. Then, at $t=0$, the system is quenched to the bath temperature $T_{\text{b}}^{+}$, so that $\lim_{t\to\infty}T(t)=T_{\text{b}}^{+}$.
@@
$T(t)=\begin{cases}T_{\text{b}}^{-},&t<0,\\ T_{\text{b}}^{+}+\left(T_{\text{b}}^{-}-T_{\text{b}}^{+}\right)\mathcal{E}(t),&t>0,\end{cases}$ (12)
$\mathcal{E}(t)$ being the inverse Laplace transform of ~$\widetilde{\mathcal{E}}(s)$. In the limit of no delay ($\tau\to 0$), ~$\widetilde{\mathcal{E}}(s)\to(1+s)^{-1}$, so that $\mathcal{E}(t)\to e^{-t}$ and Eq. (2) is recovered. Because of that, henceforth the function $\mathcal{E}(t)$ will be referred to as the $\tau$-exp function .
@@
$=$ $1+\sum_{n=0}^{\lfloor{t/\tau}\rfloor}\frac{(n\tau-t)^{n+1}}{(n+1)!},\quad t>0,$ (14)
@@
$\mathcal{E}(t)=\begin{cases}1-t,&0<t<\tau,\\ 1-t+\frac{(\tau-t)^{2}}{2},&\tau<t<2\tau,\\ 1-t+\frac{(\tau-t)^{2}}{2}+\frac{(2\tau-t)^{3}}{3!},&2\tau<t<3\tau,\\ \cdots,&\cdots\end{cases}$ (15)
@@
˙$\dot{\mathcal{E}}(t)=-\mathcal{E}(t-\tau).$ (16)
Figure 3: Plot of the $\tau$-exp function $\mathcal{E}(t)$ with, from top to bottom, $\tau=0$, $0.1$, $0.2$, $0.3$, and $e^{-1}\simeq 0.368$. The inset shows the graph in semilogarithmic scale.
Figure 3 shows the $\tau$-exp function $\mathcal{E}(t)$ for several values of $\tau$ in normal (main panel) and semilogarithmic (inset) scales. As we can see, the larger the time delay $\tau$, the faster the decay of $\mathcal{E}(t)$.
The long-time behavior of $\mathcal{E}(t)$ is governed by the dominant pole of ~$\widetilde{\mathcal{E}}(s)$, that is, the root of the transcendental equation $s+e^{-s\tau}=0$ with the least negative real part. It can be shown that such a root is real ($s=-\kappa$ with $\kappa<\tau^{-1}$) provided that $\tau<\tau_{\max}=e^{-1}\simeq 0.368$ (at which case $\kappa=e\simeq 2.718$). In that domain, $\kappa$ is the real root of $\kappa=e^{\kappa\tau}$, i.e., $\kappa=-\tau^{-1}W_{0}(-\tau)$, where $W_{0}(z)$ is the principal branch of the Lambert function. If, on the other hand, $\tau>\tau_{\max}$, then the dominant root is a pair of complex conjugates, implying an oscillatory behavior with a negative minimum $\mathcal{E}_{\min}=\mathcal{E}(t_{\min})<0$ at a certain time $t=t_{\min}$. The existence of this minimum compromises the positive-definiteness of the solution, Eq. (12). Suppose that $\tau>\tau_{\max}=e^{-1}$ and $T_{\text{b}}^{-}/T_{\text{b}}^{+}>1+|\mathcal{E}_{\min}|^{-1}$. Then, $T(t_{\min})/T_{\text{b}}^{+}=1-(T_{\text{b}}^{-}/T_{\text{b}}^{+}-1)|\mathcal{E}_{\min}|<0$. Therefore, the time-delayed Newton’s cooling law is physically meaningful only if $\tau<\tau_{\max}=e^{-1}$. This is the maximum value of $\tau$ considered in Fig. 3.
Figure 4: Plot of the damping coefficient $\kappa$ and the amplitude $A_{\mathcal{E}}$ as functions of $\tau$. The circle represents the damping coefficient $\kappa=e\simeq 2.718$ at the maximum time delay $\tau_{\max}=e^{-1}\simeq 0.368$.
Figure 5: Schematic representation of the decomposition of the double-quench protocol (a) into the sum of the single-quench protocols (b) and (c).
@@
The damping coefficient $\kappa$ and the amplitude $A_{\mathcal{E}}$ are displayed as functions of the time delay $\tau$ in Fig. 4. The damping coefficient increases from $\kappa=1$ at $\tau=0$ to $\kappa=e\simeq 2.718$ at the maximum time delay $\tau_{\max}=e^{-1}\simeq 0.368$. As for the amplitude, it is practically $A_{\mathcal{E}}\simeq 1$ until $\tau\simeq 0.3$, but then it diverges as $\tau$ approaches its maximum value.
@@
$T_{\text{A}}(t)-T_{\text{B}}(t)=\begin{cases}T_{\text{b,A}}^{-}-T_{\text{b,B}}^{-},&t<0,\\ (T_{\text{b,A}}^{-}-T_{\text{b,B}}^{-})\mathcal{E}(t),&t>0.\end{cases}$ (18)
This proves that no Mpemba effect is possible with the one-quench protocol.
II.3 A more complex solution. Double quench
Now, instead of taking the single-quench protocol, Eq. (8), we us assume that the material is kept at a prior temperature $T_{\text{b}}^{-}$ for times $t<-t_{\text{w}}$, next it is quenched to a middle bath temperature $T_{\text{b}}^{-*}$ for $-t_{w}<t<0$, and then (after a waiting time $t_{\text{w}}$) it is quenched again to a final bath temperature $T_{\text{b}}^{+}$. This double-quench protocol is
$T_{\text{b}}(t)=\begin{cases}T_{\text{b}}^{-},&t<-t_{\text{w}},\\ T_{\text{b}}^{-*},&-t_{\text{w}}<t<0.\\ T_{\text{b}}^{+},&t>0.\end{cases}$ (19)
@@
$T_{1}(t)=\begin{cases}T_{\text{b}}^{-}-T_{\text{b}}^{-*}+T_{\text{b}}^{+}/2,&t<-t_{\text{w}},\\ T_{\text{b}}^{+}/2+(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&t>-t_{\text{w}},\end{cases}$ (21a)
$T_{2}(t)=\begin{cases}T_{\text{b}}^{-*}-T_{\text{b}}^{+}/2,&t<0,\\ T_{\text{b}}^{+}/2+(T_{\text{b}}^{-*}-T_{\text{b}}^{+})\mathcal{E}(t),&t>0.\end{cases}$ (21b)

### v2

@@ Abstract
Despite extensive research, the fundamental physical mechanisms underlying the Mpemba effect, a phenomenon where a substance cools faster after initially being heated, remain elusive. Although historically linked with water, the Mpemba effect manifests across diverse systems, sparking heightened interest in Mpemba-like phenomena. Concurrently, the Kovacs effect, a memory phenomenon observed in materials such as polymers, involves rapid quenching and subsequent temperature changes, resulting in nonmonotonic relaxation behavior. This paper probes the intricacies of the Mpemba and Kovacs effects within the framework of the time-delayed Newton’s law of cooling, recognized as a simplistic yet effective phenomenological model accommodating memory phenomena. This law allows for a nuanced comprehension of temperature variations, introducing a delay time ($\tau$) and incorporating specific protocols for the thermal bath temperature, contingent on a defined waiting time ($t_{\text{w}}$). Remarkably, the relevant parameter space is two-dimensional ($\tau$ and $t_{\text{w}}$), with bath temperatures exerting no influence on the presence or absence of the Mpemba effect or the relative strength of the Kovacs effect. The findings enhance our understanding of these memory phenomena, providing valuable insights applicable to researchers across diverse fields, ranging from physics to materials science.
@@ I Introduction
As a teenager, Mpemba (1950–2023) accidentally discovered the paradoxical effect that now bears his name. In the first part of his renowned paper [1] with Osborne (1932–2014), the 19-year-old Mpemba candidly recounted the story. In the second part of the paper, Osborne wrote,
“The headmaster of Mkwawa High School invited me to speak to the students on ‘Physics and national development’. […] One student raised a laugh from his colleagues with a question I remember as ‘If you take two beakers with equal volumes of water, one at $35^{\circ}$ and the other at $100^{\circ}$, and put them into a refrigerator, the one that started at $100^{\circ}$ freezes first. Why?’. It seemed an unlikely happening, but the student insisted that he was sure of the facts. I confess that I thought he was mistaken but fortunately remembered the need to encourage students to develop questioning and critical attitudes. No question should be ridiculed. In this case there was an added reason for caution, for everyday events are seldom as simple as they seem and it is dangerous to pass a superficial judgment on what can and cannot be. I said that the facts as they were given surprised me because they appeared to contradict the physics I knew. But I added that it was possible that the rate of cooling might be affected by some factor I had not considered. I promised I would put the claim to the test of experiment and encouraged my questioner to repeat the experiment himself.”
In that second part of Ref. [1], Osborne reported some experimental results confirming the effect, although he conceded that “The experiments attempted were relatively crude and several factors could influence cooling rates. More sophisticated experiments are needed to provide a more certain answer to the question.”
@@
After 1969, attention to the Mpemba effect was mainly confined to popular science and education journals [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. It was also revealed that the phenomenon had already been noted by classical philosophers and scholars, including Aristotle [24], Roger Bacon [25], Francis Bacon [26], and Descartes [27]. Interestingly, in the same year that Mpemba and Osborne published their paper, Kell independently published a brief note that began with the following statement: “It is widely believed, at least in Canada, that hot water will freeze more quickly than cold water.”
@@
In addition to its historical association with water, the Mpemba effect exhibits similar phenomena reported across a diverse range of systems, including carbon nanotube resonators [41], clathrate hydrates [42], Markovian models [43, 44, 45, 46, 47, 48], granular gases [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], molecular gases under drag [60, 61, 62, 59], spin glasses [63], non-Markovian mean-field systems [64, 65], inertial suspensions [66, 67], colloidal systems [68, 69, 45, 70, 71], Ising-like models [72, 73, 74, 48, 75, 76], ideal gases [77], phase transitions [78], active matter [79], the autonomous information engine [80], ionic liquids [81], polymers [82, 83], Langevin systems [84, 85], and plasmas [86]. Particularly noteworthy is the recent surge in studies examining the effect in quantum systems [87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]. The escalating interest in Mpemba-like effects is illustrated by Fig. 1.
Figure 1: Number of publications and citations in the period 2006–2023 obtained from the Web of Science [98] with the search query “topic=(Mpemba effect).” Note the vertical logarithmic scale.
Figure 2: Sketch of supposed cooling curves for two similar systems in the same environment, adapted from Ref. [8].
@@
where $T_{\text{b}}$ is the temperature of the thermal bath (or environment) and $\lambda$ is the coefficient of heat transfer (or cooling rate), here assumed to be a (positive) constant. Its value depends on various factors including the material properties, the object geometry, the surface conditions, and the surrounding environment. For instance, in the case of water, $\lambda$ can typically range from about $10^{-3}~{}\text{s}$ to $10^{-2}~{}\text{s}$ depending on the temperature difference and the volume of water [102].
The general solution to Eq. (1) is simply
@@
with $T_{0}=T(0)$ being the initial temperature. On the other hand, if the Mpemba effect exists in a certain material, memory dynamics must be taken into account, so that “the overtaking effect should enable us to describe the difference between the two systems as represented by the points B and D [see Fig. 2] in such a way that we would expect the system starting hotter to cool faster even below this temperature.” The simplest way to incorporate those memory phenomena into Newton’s law involves postulating that the temperature’s rate of change at time $t$ depends on the temperature at a preceding time $t-\tau$ [103], i.e.,
@@
where $\tau>0$ is the delay time and we have considered the possibility that the bath temperature, $T_{\text{b}}(t)$, changes with time. Equation (3) can be considered as the simplest phenomenological equation incorporating memory effects.
@@
Additionally to the Mpemba effect, another fascinating memory phenomenon is the Kovacs effect, originally reported in polymer materials [104, 105] and also observed in other complex systems [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 57, 59]. In the Kovacs effect, a sample, initially in equilibrium at a high temperature $T_{\text{b}}^{-}$, undergoes rapid quenching to a lower temperature $T_{\text{b}}^{+}$. The sample evolves for a specified waiting time $t_{\text{w}}$, but subsequently the bath temperature is abruptly raised to $T(t_{\text{w}})$. Beyond $t=t_{\text{w}}$, the sample temperature exhibits a dynamic behavior: it initially increases, reaches a maximum, and then returns to equilibrium for a longer period. The effect thus highlights the nontrivial impact of the material’s thermal past on its present and future behavior.
@@
The aim of this paper is to analyze the Mpemba and Kovacs effects as described by the time-delayed cooling law, given by Eq. (3). Its general solution is studied in Sec. II, with special attention to single and double quenches. It is seen that, in the former case, the solution has a form similar to Eq. (2), except that the role of the exponential $e^{-\lambda t}$ is played by a function $\mathcal{E}(t)$, here called the $\tau$-exp function. This function is positive definite only if the delay time is smaller than a threshold value, $\tau_{\max}=\lambda^{-1}e^{-1}$. Next, in Sec. III, the solution is applied to the study of the Mpemba effect under simple protocols involving a hot bath at temperature $T_{\text{b}}^{\text{h}}$ and a cold bath at temperature $T_{\text{b}}^{\text{c}}$. Sample A is thermalized at temperature $T_{\text{b}}^{\text{h}}$ and then quenched to temperature $T_{\text{b}}^{\text{c}}$. Conversely, sample B is thermalized at $T_{\text{b}}^{\text{c}}$, quenched to $T_{\text{b}}^{\text{h}}$, and, after a waiting time $t_{\text{w}}$, quenched again to the cold temperature $T_{\text{b}}^{\text{c}}$. For simplicity, the quench of sample A and the first quench of sample B occur at the same time. After this preparation protocol, both samples are in the same environment (bath $T_{\text{b}}^{\text{c}}$) but start with different temperatures. It is proved in Sec. III that (i) the existence or absence of the Mpemba effect is independent of the values of $T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$, (ii) the Mpemba effect exists if and only if the two control parameters ($\tau$ and $t_{\text{w}}$) lie inside a certain narrow region, and (iii) the direct (cooling process) and the inverse (heating process) effects are fully equivalent. Section IV is devoted to the Kovacs effect and the function characterizing the strength of the Kovacs hump is identified. Interestingly, this hump points downward in the cooling process, thus signaling the presence of an anomalous Kovacs effect [108, 114], which becomes relatively stronger as both the delay time $\tau$ and the waiting time $t_{\text{w}}$ increase. Finally, the paper is closed in Sec. V with a summary and conclusion.
@@ II Time-delayed Newton’s cooling law
II.1 Expansion in powers of the delay time
Before proceeding with the complete delayed equation, given by Eq. (3), let us consider $\tau$ as a small parameter and expand $T(t-\tau)$ in a series of powers of $\tau$:
$T(t-\tau)=\sum_{k=0}^{\infty}\frac{d^{k}T(t)}{dt^{k}}\frac{(-\tau)^{k}}{k!}.$ (4)
Truncation at the level $k=0$ yields the original law, given by Eq. (1). Interestingly, truncation at $k=1$ just produces the same law, except for a renormalized cooling rate $\lambda^{\prime}\equiv\lambda/(1-\lambda\tau)$. However, the equation resulting from truncation at $k=2$ yields
$\mu\ddot{T}(t)+\dot{T}(t)=-\lambda^{\prime}\left[T(t)-T_{\text{b}}(t)\right],\quad\mu\equiv\frac{\lambda^{\prime}\tau^{2}}{2}.$ (5)
The presence of the term $\mu\ddot{T}(t)$ suggests a more complex heat transfer process compared to the standard Newton’s cooling equation. It represents a sort of “thermal inertia” of the material, meaning that it might resist changes in temperature more than as described by Newton’s law. In fact, by assuming $T_{\text{b}}=\text{const}$, the general solution to Eq. (5) is
$T(t)=$ $T_{\text{b}}+\frac{T_{0}-T_{\text{b}}}{\lambda_{+}^{\prime}-\lambda_{-}^{\prime}}\left(\lambda_{+}^{\prime}e^{-\lambda_{-}^{\prime}t}-\lambda_{-}^{\prime}e^{-\lambda_{+}^{\prime}t}\right)$
$+\frac{\dot{T}_{0}}{\lambda_{+}^{\prime}-\lambda_{-}^{\prime}}\left(e^{-\lambda_{-}^{\prime}t}-e^{-\lambda_{+}^{\prime}t}\right),$ (6)
where
$\lambda_{\pm}^{\prime}\equiv\frac{1\pm\sqrt{1-4\lambda^{\prime}\mu}}{2\mu},\quad\dot{T}_{0}\equiv\dot{T}(0).$ (7)
Thus, the temperature evolution is not solely determined by the initial temperature $T_{0}$, as is the case with Eq. (2), but also by the initial slope $\dot{T}_{0}$, thereby violating the simplistic depiction outlined in Fig. 2 and allowing memory effects to manifest. It is worthwhile noting that the second-order differential equation, given by Eq. (5), with $T_{\text{b}}=\text{const}$ is equivalent to the following coupled set of two first-order differential equations:
$\dot{T}(t)=-\lambda^{\prime}\left[T(t)-T_{\text{b}}\right]+q(t),$ (8a)
$\dot{q}(t)=-{\lambda^{\prime}}^{2}\left[T(t)-T_{\text{b}}\right]-\left(\mu^{-1}-\lambda^{\prime}\right)q(t).$ (8b)
This two-variable scheme resembles the approach used to investigate the Mpemba and Kovacs effects in granular and molecular gases [49, 50, 51, 60, 66, 53, 55, 56, 62, 58].
Certainly, truncating Eq. (4) at higher orders leads to differential equations of increasing order, thereby amplifying the influence of memory effects.
II.2 General solution in Laplace space
The general solution of Eq. (3) depends not only on the initial temperature $T_{0}$ but it is actually a functional of the previous history $T(t)$ for $-\tau<t<0$. Given the linear character of Eq. (3), it is convenient to define the Laplace transforms as
@@ II.2 General solution in Laplace space
and, for simplicity, we have taken $\lambda^{-1}=1$ as the unit of time. Thus, the general solution in Laplace space is
@@
II.3 Basic solution. Single quench
@@
The system is assumed to have equilibrated at $T_{\text{b}}^{-}$ for $t<0$, that is, $T(t\leq 0)=T_{\text{b}}^{-}$. Then, at $t=0$, the system is quenched to the bath temperature $T_{\text{b}}^{+}$, so that $\lim_{t\to\infty}T(t)=T_{\text{b}}^{+}$.
@@
$T(t)=\begin{cases}T_{\text{b}}^{-},&t\leq 0,\\ T_{\text{b}}^{+}+\left(T_{\text{b}}^{-}-T_{\text{b}}^{+}\right)\mathcal{E}(t),&t\geq 0,\end{cases}$ (17)
with $\mathcal{E}(t)$ being the inverse Laplace transform of ~$\widetilde{\mathcal{E}}(s)$. In the limit of no delay ($\tau\to 0$), ~$\widetilde{\mathcal{E}}(s)\to(1+s)^{-1}$, so that $\mathcal{E}(t)\to e^{-t}$ and Eq. (2) is recovered. Because of that, henceforth the quasi-exponential function $\mathcal{E}(t)$ will be referred to as the $\tau$-exp function .
@@
$=$ $1+\sum_{n=0}^{\lfloor{t/\tau}\rfloor}\frac{(n\tau-t)^{n+1}}{(n+1)!},\quad t\geq 0,$ (19)
@@
$\mathcal{E}(t)=\begin{cases}1-t,&0\leq t\leq\tau,\\ 1-t+\frac{(\tau-t)^{2}}{2},&\tau\leq t\leq 2\tau,\\ 1-t+\frac{(\tau-t)^{2}}{2}+\frac{(2\tau-t)^{3}}{3!},&2\tau\leq t\leq 3\tau,\\ \cdots,&\cdots\end{cases}$ (20)
@@
˙$\dot{\mathcal{E}}(t)=-\begin{cases}1,&0\leq t\leq\tau,\\ \mathcal{E}(t-\tau),&t\geq\tau.\end{cases}$ (21)
Figure 3: Plot of the $\tau$-exp function $\mathcal{E}(t)$ with, from top to bottom, $\tau=0$, $0.1$, $0.2$, $0.3$, and $e^{-1}\simeq 0.368$. Inset; The graph in semilogarithmic scale.
Figure 3 shows the $\tau$-exp function $\mathcal{E}(t)$ for several values of $\tau$ in normal (main panel) and semilogarithmic (inset) scales. As we can see, the larger the delay time $\tau$, the faster the decay of $\mathcal{E}(t)$.
The long-time behavior of $\mathcal{E}(t)$ is governed by the dominant pole of ~$\widetilde{\mathcal{E}}(s)$, that is, the root of the transcendental equation $s+e^{-s\tau}=0$ with the least negative real part. It can be shown that such a root is real ($s=-\kappa$ with $\kappa<\tau^{-1}$) provided that $\tau<\tau_{\max}=e^{-1}\simeq 0.368$ (at which case $\kappa=e\simeq 2.718$). In that domain, $\kappa$ is the real root of $\kappa=e^{\kappa\tau}$, i.e., $\kappa=-\tau^{-1}W_{0}(-\tau)$, where $W_{0}(z)$ is the principal branch of the Lambert function. If, on the other hand, $\tau>\tau_{\max}$, then the dominant root is a pair of complex conjugates, implying an oscillatory behavior with a negative absolute minimum $\mathcal{E}_{\min}=\mathcal{E}(t_{\min})<0$ at a certain time $t=t_{\min}$. The existence of this minimum compromises the positive-definiteness of the solution, given by Eq. (17). Suppose that $\tau>\tau_{\max}=e^{-1}$ and $T_{\text{b}}^{-}/T_{\text{b}}^{+}>1+|\mathcal{E}_{\min}|^{-1}$. Then, $T(t_{\min})/T_{\text{b}}^{+}=1-(T_{\text{b}}^{-}/T_{\text{b}}^{+}-1)|\mathcal{E}_{\min}|<0$. Therefore, the time-delayed Newton’s cooling law is physically meaningful only if $\tau<\tau_{\max}=e^{-1}$. This is the maximum value of $\tau$ considered in Fig. 3.
Figure 4: Plot of the damping coefficient $\kappa$ and the amplitude $A_{\mathcal{E}}$ as functions of $\tau$. The circle represents the damping coefficient $\kappa=e\simeq 2.718$ at the maximum delay time $\tau_{\max}=e^{-1}\simeq 0.368$.
Figure 5: Schematic representation of the decomposition of the (a) double-quench protocol into the sum of the (b),(c) single-quench protocols.
@@
The damping coefficient $\kappa$ and the amplitude $A_{\mathcal{E}}$ are displayed as functions of the delay time $\tau$ in Fig. 4. The damping coefficient increases from $\kappa=1$ at $\tau=0$ to $\kappa=e\simeq 2.718$ at the maximum delay time $\tau_{\max}=e^{-1}\simeq 0.368$. As for the amplitude, it is practically $A_{\mathcal{E}}\simeq 1$ until $\tau\simeq 0.3$, but then it diverges as $\tau$ approaches its maximum value.
@@
$T_{\text{A}}(t)-T_{\text{B}}(t)=\begin{cases}T_{\text{b,A}}^{-}-T_{\text{b,B}}^{-},&t\leq 0,\\ (T_{\text{b,A}}^{-}-T_{\text{b,B}}^{-})\mathcal{E}(t),&t\geq 0.\end{cases}$ (23)
Since $\mathcal{E}(t)>0$ if $\tau<\tau_{\text{max}}$, Eq. (23) proves that no Mpemba effect is possible with the single-quench protocol.
II.4 A more complex solution. Double quench
Now, instead of taking the single-quench protocol, given by Eq. (13), let us assume that the material is kept at a prior temperature $T_{\text{b}}^{-}$ for times $t<-t_{\text{w}}$, next it is quenched to a middle bath temperature $T_{\text{b}}^{-*}$ for $-t_{w}<t<0$, and then (after a waiting time $t_{\text{w}}$) it is quenched again to a final bath temperature $T_{\text{b}}^{+}$. This double-quench protocol is
$T_{\text{b}}(t)=\begin{cases}T_{\text{b}}^{-},&t<-t_{\text{w}},\\ T_{\text{b}}^{-*},&-t_{\text{w}}<t<0,\\ T_{\text{b}}^{+},&t>0.\end{cases}$ (24)
@@
$T_{1}(t)=\begin{cases}T_{\text{b}}^{-}-T_{\text{b}}^{-*}+T_{\text{b}}^{+}/2,&t\leq-t_{\text{w}},\\ T_{\text{b}}^{+}/2+(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&t\geq-t_{\text{w}},\end{cases}$ (26a)